### v1

# Advancing Universal Deep Learning for Electronic-Structure Hamiltonian Prediction of Materials

Abstract
[…]
1 Introduction
Understanding the electronic structure is fundamental to unraveling how electrons govern the properties of condensed matter systems. This knowledge is essential for predicting a wide range of material characteristics, such as electrical conductivity, magnetism, optical behavior, and chemical activity, which are vital for technologies spanning from electronics to sustainable energy and advanced catalysis. At the heart of these calculations is the challenge of determining the system’s Hamiltonian matrix, whose eigenvalues and eigenstates yield important quantities like energy levels, band structures, and electronic wavefunctions. Traditionally, Density Functional Theory (DFT) (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) has been the standard approach for these problems. However, as shown in Fig. 1 (a), DFT relies heavily on the self-consistent (SC) procedure, which demands repeated (denoted as $T$ turns), computationally intensive diagonalizations of large matrices—each scaling as $\mathcal{O}(N^{3})$ with system size $N$, making simulations of large or complex materials extremely resource-consuming. Recently, deep learning has emerged as a powerful tool in the physical sciences (Zhang et al., 2025). As shown in Fig. 1 (b), modern deep neural network methods (Gong et al., 2023; Yu et al., 2023; Zhang et al., 2024; Wang et al., 2024b; Li et al., 2025; Yin et al., 2025) can predict Hamiltonians, i.e., the core physical quantities in electronic structure calculations, directly from atomic configurations in an efficient way, circumventing the computationally expensive SC loop and dramatically accelerating computations. This paradigm shift lowers the computational barriers associated with electronic structure calculations, unlocking the simulation and design of unprecedentedly large-scale materials systems, driving new innovation in materials discovery and engineering. Please refer to Appendix A.1 for background introduction.
Figure 1: Comparison of paradigms for electronic-structure Hamiltonian calculation, highlighting the fundamental differences between our method and both classical DFT methods and existing deep learning approaches.
However, deep learning methods still face substantial challenges in achieving accurate and generalizable Hamiltonian prediction, primarily due to the extremely complex and fundamentally difficult nature of the input–output mapping that the neural network must learn, making it difficult to generalize across diverse material systems. Consequently, it has become common practice to constrain the scope, such as limiting the range of supported elements, neglecting spin–orbit coupling (SOC) effects, or reducing the number of orbitals considered, as discussed in Section 2. While such strategies help alleviate modeling burdens, they also restrict the applicability of these methods to the full diversity and complexity of real-world materials. What’s more, large open-source materials datasets for the training and evaluation of general Hamiltonian learning models are also rare.
To solve these challenges, in this work, we make contributions on both methodology and benchmark toward advancing universal deep learning for electronic-structure Hamiltonian prediction of materials. On the method side, as shown in Fig. 1 (c), we propose NextHAM, a neural E(3)-symmetry and expressive correction framework for efficient and accurate Hamiltonian prediction:
First, we dive deeply into the traditional DFT computational process outlined in Appendix A.1 and novely introduce a physical quantity that helps mitigate the complexity of the input–output mapping encountered by deep learning models for Hamiltonian prediction. This quantity is the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$, which is efficiently constructed from the initial electron density $\rho^{(0)}(\mathbf{r})$, given by the sum of the charge densities of isolated atoms, without the requirement of matrix diagonalization. As $\mathbf{H}^{(0)}$ efficiently encodes essential information about the system’s electronic structure, we use it as one of the input features to the neural network. Unlike existing methods that rely on randomly initialized atom and edge embeddings lacking physical prior knowledge, $\mathbf{H}^{(0)}$ provides richer physical context by embedding the intrinsic characteristics of diverse elements into a unified representation space, thereby enabling robust generalization across chemically complex material systems. Furthermore, we make our method a correction approach by taking $\Delta\mathbf{H}=\mathbf{H}^{(T)}-\mathbf{H}^{(0)}$ instead of $\mathbf{H}^{(T)}$ as the regression target of neural networks, significantly reducing their output space. This compression of the output space helps the regression model to finely fit and generalize.
Second, we present a network architecture that strictly adheres to E(3) symmetry while maintaining high non-linear expressiveness for Hamiltonian prediction by extending the TraceGrad (Yin et al., 2025) method to Transformer framework, thereby providing ample capacity for flexible and accurate modeling of atomic systems for Hamiltonian prediction across a wide range of elements in the periodic table. Furthermore, we introduce model ensemble techniques to enhance the capacity of the framework for handling complex scenarios in Hamiltonian prediction.
Third, we propose a joint optimization framework that simultaneously refines both real-space (R-space) and reciprocal-space (k-space) Hamiltonians. Most existing methods regress only the real-space Hamiltonian, but the large condition number of the overlap matrix can amplify errors in predicted eigenvalues and eigenfunctions, leading to suboptimal physical fidelity. By jointly optimizing R- and k-space Hamiltonians with a unified loss, our approach ensures accurate prediction of key electronic properties, e.g. the band structures.
On the dataset side, we curate a diverse-collection large benchmark dataset, Materials-HAM-SOC, containing $17,000$ material structures generated using DFT softwares. The dataset spans $68$ elements from the first six rows of the periodic table and explicitly incorporates spin–orbit coupling (SOC) effects. To ensure the accuracy of the DFT calculations, we employ high-quality pseudopotentials that include as many valence electrons as possible, enabling our model to handle physically complex and highly challenging systems. We adopt high-quality atomic orbital basis sets, up to 4s2p2d1f orbitals for each element, to ensure fine-grained description of electronic structures. This dataset establishes a challenging yet comprehensive benchmark for evaluating generalization across chemically and structurally diverse systems.
[…]
2 Related Work
The foundation of deep learning-based electronic-structure Hamiltonian prediction lies in constructing neural networks that respect E(3)-symmetry (with SO(3) as its rotational sub-group), which can be achieved by incorporating group theory-based tensor operators (Geiger & Smidt, 2022). Since traditional non-linear activation functions, when applied directly to SO(3)-equivariant features, may break equivariance, a central research topic is to reconcile strong non-linear expressiveness with strict SO(3)-equivariance. An early attempt to address this problem was the use of gated activation functions (Weiler et al., 2018), which first apply non-linear activations to SO(3)-invariant features and then use them as coefficients to scale the SO(3)-equivariant features. Representative works adopting this mechanism include Equiformer (Liao & Smidt, 2023), eSEN (Fu et al., 2025) for force prediction, and DeepH-E3 (Gong et al., 2023), QHNet (Yu et al., 2023) for Hamiltonian prediction. To further enhance non-linear expressiveness of E(3)-networks, Zitnick et al. (2022) and Passaro & Zitnick (2023) proposed eSCN (equivariant Spherical Channel Networks), which applies non-linear operations to the coefficients obtained from the spherical decomposition of features. This approach has been widely used in tasks such as force-field prediction (Liao et al., 2024) and Hamiltonian prediction (Wang et al., 2024a; b). Nevertheless, eSCN methods project features onto discrete basis functions via inner-product operations, degrading strict SO(3)-equivariance to equivariance with respect to only a discrete sub-group. More recently, Yin et al. (2025) proposed the TraceGrad method, which introduces an SO(3)-invariant supervision signal, i.e., the trace quantity, constructed from Hamiltonian labels, to supervise high-fidelity non-linear SO(3)-invariant features. Then, non-linearity is delivered to SO(3)-equivariant features via a gradient-based mechanism from the SO(3)-invariant ones. This method effectively unifies strict SO(3)-equivariance with strong non-linear expressiveness for Hamiltonian prediction; however, the backbone network it adopted is a simple graph neural network and has not yet evolved into a non-linear equivariant Transformer framework.
Despite these progresses, deep learning methods for Hamiltonian prediction still face substantial challenges on generalization performance, which can be summarized as follows. First, crystalline materials commonly found in nature can be composed of over $65$ different elements from the first six rows of the periodic table, leading to an exceptionally large and heterogeneous input space for deep neural network models. Existing deep learning methods for Hamiltonian prediction typically employ learnable embedding techniques similar to word vectors (Mikolov et al., 2013) to represent nodes (atoms) and edges (atom pairs). These embeddings are randomly initialized and learned directly from the dataset, without incorporating any explicit physical priors. As a result, they struggle to capture the fundamental physical relationships between different atoms and across different material systems, which are crucial for generalization. Second, as illustrated in Figure 1, the regression target, namely the self-consistent electronic-structure Hamiltonian, is inherently high-dimensional and complex, especially when considering SOC effects. For instance, a system containing several tens of atoms may involve nearly several thousands of non-zero Hamiltonian matrix elements that need to be accurately predicted. Most of the existing methods attempt to directly predict the entire self-consistent Hamiltonian matrix, namely $\mathbf{H}^{(T)}$ as formulated in Appendix A.1, placing a heavy burden on the model due to the vast size of the output space, often resulting in optimization difficulties during training and limited generalization to unseen systems. In addition, most existing methods treat the real-space Hamiltonian as the sole regression target, which can lead to sub-optimal physical fidelity in down-stream applications, particularly in capturing low-energy band structures accurately. Although Li et al. (2025) designed a method for molecular systems to reduce the regression space of Hamiltonians and introduced a basis transformation of the Hamiltonian matrix on the loss function to improve the prediction accuracy of down-stream physical quantities, their approach is limited to molecular systems and is not applicable to periodic crystalline materials.
As a result, constructing a unified model that generalizes across diverse crystal prototypes remains challenging, and many existing approaches explicitly constrain their scope. For example, Li et al. (2022), Gong et al. (2023), and Xia et al. (2025) each train and evaluate their methods within a single material system (e.g., MoS2, Bi2Se3, or a-HfO2), without assessing cross-material generalization. More recently, DeepH-2 (Wang et al., 2024b) broadened coverage to systems involving elements primarily from the first four rows of the periodic table; however, they neglected SOC and reduced the orbital basis by omitting $f$-orbitals. While such choices help reduce computational and modeling complexity, they may limit broad applicability to the full diversity of real-world materials. Zhong et al. (2025) developed a Hamiltonian prediction model aimed at a broader range of element types, while also highlighting the challenge of achieving consistently high accuracy across diverse crystal systems. Moreover, open-source datasets with a broad and diverse collection of materials dedicated to training and validating universal Hamiltonian models across the periodic table remain scarce.
To solve these challenges, this work presents an advanced unified deep learning framework together with a large benchmark dataset for Hamiltonian prediction, targeting broader generalization across richer classes of materials.
Figure 2: Illustration of the proposed NextHAM framework.
3 Method
To handle wider chemical and structural variability while maintaining accuracy and efficiency, we develop an unified Hamiltonian prediction framework along three fronts: (i) input descriptors and output targets, (ii) network architectures, and (iii) training objectives.
3.1 Input Descriptors and Output Targets
As shown at the lower part of Fig. 2, we use the displacement vector-based descriptors between atoms that lie within the cutoff distance, together with the zeroth-step Hamiltonian (calculated as detailed in Appendix A.1), as the input features for the neural network. Introducing the zeroth-step Hamiltonian as the input features for the neural network is one of the core innovations of our framework. The zeroth-step Hamiltonian $\mathbf{H}^{(0)}$, derived from the initial charge density $\rho^{(0)}(\mathbf{r})$, obtained as the sum of neutral atomic charges, reflects the information of different elements in the system, including the strength of the electron-ion interactions (pseudopotential) and a preliminary estimate of the electron-electron interactions. These components directly influence the system’s electronic structure. This method encodes the characteristics of different elements into a unified representation space, which enhances the model’s ability to generalize across diverse material systems.
The zeroth-step Hamiltonian $\mathbf{H}^{(0)}$ can be decomposed into on-site sub-matrices, which represent the Hamiltonian blocks corresponding to each atom and its own orbitals, and off-site sub-matrices, which capture the interactions between different atoms. These two types of sub-matrices can naturally serve as the initial descriptors for nodes and edges, respectively, in the graph neural network.
Moreover, as detailed in Appendix A.1, computing the zeroth-step Hamiltonian requires no matrix diagonalization, so its cost scales with the number of non-zero matrix elements: it is approximately $\mathcal{O}(N^{2})$ for small systems with $N$ atoms and asymptotically approaches $\mathcal{O}(N)$ for sufficiently large systems as the neighbor count saturates for each atom. This matches the scaling behavior of message passing mechanism of graph neural networks, ensuring that incorporating the zeroth-step Hamiltonian as a new input descriptor does not worsen the $\mathcal{O}$-asymptotics.
On the output side of the neural network, our approach predicts the correction term $\Delta\mathbf{H}=\mathbf{H}^{(T)}-\mathbf{H}^{(0)}$. This residual formulation reframes the learning task as a refinement problem, where the network adjusts the physically informed zeroth-step Hamiltonian $\mathbf{H}^{(0)}$ toward its self-consistent counterpart. By significantly reducing both the dimensionality and numerical range of the regression target, this approach simplifies the learning problem and enables the model to focus on capturing only the essential differences rather than reconstructing the entire Hamiltonian from scratch. As a result, it promotes generalization performance across diverse atomic configurations.
3.2 Neural Network Architecture
We present a Transformer architecture that not only maintains strict E(3)-symmetry but also achieves strong non-linear expressiveness, as shown in Fig. 2. Our E(3)-symmetry graph attention mechanism is developed from Equiformer (Liao & Smidt, 2023). While Equiformer was designed for regression tasks where the target quantity is essentially a node-level atomic property (e.g., force fields), our Hamiltonian target is fundamentally an edge-level property defined on atomic pairs. This distinction necessitates stronger modeling of edge features and motivates the development of our attention mechanism. First, we explicitly maintain and update edge features across multiple layers, rather than generating them only temporarily from node features on demand (Liao & Smidt, 2023). In this way, the computation of attention weights incorporates both the node features and the persistently maintained edge features. Second, motivated by the decay behavior of Hamiltonian matrix elements with respect to interatomic distance, we explicitly incorporate interatomic distances by introducing distance embeddings as additional signals in the computation of attention weights, enabling the model to better exploit distance information for inference. Third, the attention weights between nodes are directly applied to update edge features via multiplicative operations, and are subsequently refined through equivariant transformations. Together, these developments substantially enhance the capacity of the model to represent edge features, from which the Hamiltonian is regressed.
As analyzed in Section 2, the TraceGrad method (Yin et al., 2025) can maintain strong non-linear expressiveness while preserving strict E(3) symmetry. We extends TraceGrad into the Transformer framework for electronic-structure Hamiltonian prediction. As shown in the middle of Fig. 2, for an atomic pair $(a,b)$, the SO(3)-equivariant edge feature $\mathbf{f}^{{}^{\prime}\text{(edge)}}_{ab}$ output by the E(3)-symmetry multi-head graph attention module is fed into the TraceGrad module to produce the non-linear SO(3)-invariant feature $z^{(\text{edge})}_{ab}$, which is subsequently passed to the SO(3)-invariant decoder and trained under the supervision of the SO(3)-invariant trace quantity $\mathbf{T}=\mathrm{tr}(\Delta\mathbf{H}\cdot\Delta\mathbf{H}^{\dagger})$. The learned non-linear expressiveness in $z^{(\text{edge})}_{ab}$ is subsequently delivered into the equivariant feature by $\mathbf{o}^{\text{(edge)}}_{ab}=\mathbf{f}^{{}^{\prime}\text{(edge)}}_{ab}+\frac{\partial\,z^{(\text{edge})}_{ab}}{\partial\,\mathbf{f}^{{}^{\prime}\text{(edge)}}_{ab}}$, where $\mathbf{o}^{\text{(edge)}}_{ab}$ represents the non-linearity-enhanced SO(3)-equivariant edge feature, which, together with the node feature, are fed into the subsequent encoding modules of the Transformer as well as the SO(3)-equivariant decoder to regress the correction term $\Delta\mathbf{H}$.
To enhance model capacity and better capture the complex dependence of Hamiltonian matrix elements on diverse inter-atomic distances, we employ an ensemble learning strategy. Specifically, sub-models are trained to predict Hamiltonian sub-matrices corresponding to different distance intervals between atoms. Although each sub-model specializes in a specific range in the output stage, the input to each sub-model is the entire system, including the zeroth-step Hamiltonian and the displacement vectors for all atomic pairs, thereby effectively extracting global information. The final prediction is obtained by aggregating the outputs from all these sub-models.
3.3 Training Loss Functions
The objective of training the neural network is to make the predicted Hamiltonian, denoted as $\widehat{\mathbf{H}}=\mathbf{H}^{(0)}+\widehat{\Delta\mathbf{H}}$, approximate the ground truth $\Delta\mathbf{H}^{gt}$ as closely as possible. As illustrated in Fig. 2, to ensure that the predicted Hamiltonian can accurately derive down-stream physical quantities (such as band structures), we design a joint optimization strategy in both real space (R-space) and reciprocal space (k-space) for the neural network.
In R-space, the Hamiltonian and the corresponding trace quantity are jointly supervised. As outlined in Section 3.2, the trace quantity is used to supervise the non-linear SO(3)-invariant features, which contribute to constructing the non-linear SO(3)-equivariant features required for predicting the Hamiltonian. The R-space training loss function is defined as:
$\mathrm{loss}(\mathbf{R})={\mathbb{E}}_{\mathbf{R}}\!\left[\lambda_{R}\Big((1-\lambda_{C})\cdot\mathrm{loss}_{H}(\mathbf{R})+\gamma(\mathrm{loss}_{H},\mathrm{loss}_{T},\lambda_{C})\cdot\mathrm{loss}_{T}(\mathbf{R})\Big)\right]$ (1)
where ${\mathbb{E}}_{\mathbf{R}}[\cdot]$ denotes the empirical expectation, $\lambda_{C}$, $\lambda_{R}$ are hyperparameters; $\mathbf{R}$ denotes the lattice vector connecting the reference unit cell and a neighboring unit cell; $\mathrm{loss}_{H}(\mathbf{R})$ and $\mathrm{loss}_{T}(\mathbf{R})$ denote the prediction losses of the Hamiltonian and the trace quantity in $\mathbf{R}$-space, respectively; and $\gamma(\mathrm{loss}_{H},\mathrm{loss}_{T},\lambda_{C})$ is a scaling factor designed to balance their relative contributions for stable training. The detailed forms of these terms are provided in Appendix A.3.
As analyzed in Appendix A.2, due to the error amplification mechanism associated with the ill-conditioned overlap matrix, even small numerical errors in $R$-space can be magnified in $k$-space, leading to deviations in downstream physical quantities. To mitigate this, we introduce $k$-space loss functions. Specifically, the spectrum is partitioned into a low-energy subspace $P$, which governs most physical properties, and a high-energy complement $Q$. While downstream phenomena are predominantly determined by $P$, an inaccurately predicted Hamiltonian may introduce spurious couplings between $P$ and $Q$. This can result in unphysical abrupt changes in band structures, which are referred to as “ghost states” (see Fig. 14 in Appendix A.8). Therefore, it is essential not only to emphasize accuracy in $P$ but also to maintain reasonable fidelity in $Q$ so that the erroneous $PQ$ couplings can be identified and suppressed. To this end, we incorporate differentiated weights for $P$ and $Q$ in the loss design, together with an explicit $PQ$ penalty that eliminates unphysical cross-subspace couplings and suppresses ghost states.
The loss function in reciprocal space is defined as:
$\mathrm{loss}(\mathbf{k})={\mathbb{E}}_{\mathbf{k}}[\lambda_{P}\cdot\mathrm{loss}_{P}(\mathbf{k})+\lambda_{Q}\cdot\mathrm{loss}_{Q}(\mathbf{k})+\lambda_{PQ}\cdot\mathrm{loss}_{PQ}(\mathbf{k})]$ (2)
where $\lambda_{P}$, $\lambda_{Q}$, and $\lambda_{PQ}$ are tunable hyperparameters that adjust the relative importance of the three loss terms, which respectively measure the errors in the $P$ subspace, the $Q$ subspace, and the combined $PQ$ joint subspace. The detailed formulations of these terms are provided in Appendix A.3.
The overall loss function combines the losses from both R-space and k-space:
$\begin{split}\mathrm{loss}_{all}=\mathrm{loss}(\mathbf{R})+\mathrm{loss}(\mathbf{k})\end{split}$ (3)
This consistent treatment of real-space and reciprocal space Hamiltonians provides a robust foundation for high-fidelity band structure predictions and, in particular, effectively eliminates ghost states.
4 Dataset
As broad-coverage open-source Hamiltonian datasets that use fine-grained orbital descriptions and include spin-orbit coupling (SOC) effects across a wide range of crystals are still rare, we construct one ourselves and contribute it to the community. Specifically, our dataset, called as Materials-HAM-SOC, contains 17,00017{,}000 material structures sampled from the Materials Project (Jain et al., 2013), with ground-truth Hamiltonians and band structures generated using the DFT software ABACUS (Li et al., 2016; Lin et al., 2023) and PYATB (Jin et al., 2023). It spans $68$ distinct elements from the first six rows of the periodic table and explicitly incorporates SOC effects. For these structures, a high-quality atomic orbital basis set (Lin et al., 2021), up to 4s2p2d1f orbitals for each element, is employed, providing a fine-grained representation of their electronic structure. The dataset contains all quantities required by our method, including atomic structures, zeroth-step Hamiltonians, self-consistent Hamiltonians, and overlap matrices. The dataset is partitioned into 12,000 structures for training, 2,000 for validation, and 3,000 for testing. For details of the dataset construction and comprehensive statistical summaries, please refer to Section A.4.
5 Empirical Study
We perform empirical studies on the Materials-HAM-SOC dataset. The implementation details of the network architecture and training configurations are provided in Appendix A.5.
First, to evaluate the role of $\mathbf{H}^{(0)}$ as an initial approximation at the output stage, we measure its discrepancy from the ground truth Hamiltonian $\mathbf{H}^{gt}=\mathbf{H}^{(T)}$. This quantifies how much $\mathbf{H}^{(0)}$ reduces the effective size and complexity of the regression target space for subsequent corrections. Second, we examine the final prediction accuracy by comparing $\mathbf{H}^{(0)}+\widehat{\Delta\mathbf{H}}$ with $\mathbf{H}^{(T)}$, thereby measuring the contribution of the learned correction $\widehat{\Delta\mathbf{H}}$ in closing the residual gap between $\mathbf{H}^{(0)}$ and $\mathbf{H}^{(T)}$. These two comparisons together disentangle the effectiveness of the prior $\mathbf{H}^{(0)}$ and the neural correction on achieving high-fidelity Hamiltonian predictions.
While mean absolute error (MAE) is a straightforward error metric, Hamiltonian prediction presents a unique gauge freedom: adding a global shift $\mu\mathbf{S}$, where $\mu$ is an arbitrary scalar and $\mathbf{S}$ is the overlap matrix, leaves all down-stream physical quantities unchanged (Wang et al., 2024b). This necessitates a gauge-invariant error metric for fair evaluation. To remove this gauge freedom, we adopt the Gauge MAE (Wang et al., 2024b) to our context:
$\mathrm{Gauge\_MAE}(\mathbf{H}^{(0)},\mathbf{H}^{(T)})=\min_{\mu}\,\mathrm{MAE}\big(\mathbf{H}^{(0)},\,\mathbf{H}^{(T)}+\mu\mathbf{S}\big),$ (4)
$\mathrm{Gauge\_MAE}(\mathbf{H}^{(0)}+\widehat{\Delta\mathbf{H}},\mathbf{H}^{(T)})=\min_{\mu}\,\mathrm{MAE}\big(\mathbf{H}^{(0)}+\widehat{\Delta\mathbf{H}},\,\mathbf{H}^{(T)}+\mu\mathbf{S}\big),$
where $\mu$ is determined by solving $\mu^{*}=\operatorname*{arg\,min}_{\mu}\mathrm{Gauge\_MAE}$.
The experimental results for the above metrics are reported in Table 1. In addition, we also report $\mathrm{Gauge\_MAE}(\mathbf{0},\mathbf{H}^{(T)})$ for comparison. Comparing between $\mathrm{Gauge\_MAE}(\mathbf{0},\mathbf{H}^{(T)})$ and $\mathrm{Gauge\_MAE}(\mathbf{H}^{(0)},\mathbf{H}^{(T)})$ quantifies the actual reduction in the effective output space achieved by introducing $\mathbf{H}^{(0)}$ at the output stage.
Table 1: Comparison of Gauge MAE values computed in real space (R-space) on the testing set of Materials-HAM-SOC. Values are reported for four spin-resolved regions ($\uparrow\uparrow$, $\uparrow\downarrow$, $\downarrow\uparrow$, $\downarrow\downarrow$) with separate real and imaginary components, and for the entire matrix (Overall), where real and imaginary components are combined into a single metric. Metrics are averaged over non-zero elements only; entries set to zero due to the truncation distance are masked out. All values are in meV.
Region $\mathrm{Gauge\_MAE}(\mathbf{0},\mathbf{H}^{(T)})$ $\mathrm{Gauge\_MAE}(\mathbf{H}^{(0)},\mathbf{H}^{(T)})$ $\mathrm{Gauge\_MAE}(\mathbf{H}^{(0)}+\widehat{\Delta\mathbf{H}},\mathbf{H}^{(T)})$
Real Imag Real Imag Real Imag
$\uparrow\uparrow$ $149.145$ $0.293$ $5.213$ <0.001<0.001 $2.834$ <0.001<0.001
$\uparrow\downarrow$ $0.301$ $0.299$ <0.001<0.001 <0.001<0.001 <0.001<0.001 <0.001<0.001
$\downarrow\uparrow$ $0.301$ $0.299$ <0.001<0.001 <0.001<0.001 <0.001<0.001 <0.001<0.001
$\downarrow\downarrow$ $149.145$ $0.293$ $5.213$ <0.001<0.001 $2.834$ <0.001<0.001
Overall 74.914 2.606 1.417
As shown in Table 1, the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$ closely matches the self-consistent Hamiltonian $\mathbf{H}^{(T)}$ in the spin-flip submatrices ($\uparrow\downarrow$ and $\downarrow\uparrow$). Similarly, the imaginary parts of the spin-conserving submatrices ($\uparrow\uparrow$ and $\downarrow\downarrow$) also exhibit excellent agreement. In these components, the deviation between $\mathbf{H}^{(0)}$ and $\mathbf{H}^{(T)}$ is negligible, with errors reaching sub-$\mu eV$ level. Furthermore, in systems with time-reversal symmetry and real-valued atomic orbitals, which constitute the majority of practical cases, the real parts of the $\uparrow\uparrow$ and $\downarrow\downarrow$ blocks are identical. This symmetry implies that the correction network only needs to predict the real part of the $\uparrow\uparrow$ block in $\Delta\mathbf{H}$, substantially reducing the number of matrix elements to be learned. The $\mathrm{Gauge\_MAE}(\mathbf{H}^{(0)},\mathbf{H}^{(T)})$ for the real part of the $\uparrow\uparrow$ block is reduced by $96\%$ compared to $\mathrm{Gauge\_MAE}(\mathbf{0},\mathbf{H}^{(T)})$, yielding a much narrower numerical range for regression. This substantial reduction eases optimization by allowing the network to concentrate on physically meaningful residual corrections rather than reconstructing the entire Hamiltonian, thereby improving prediction accuracy across diverse atomic configurations. Finally, with the neural network correction applied, the errors for the $\uparrow\uparrow$ and $\downarrow\downarrow$ blocks are substantially reduced, achieving an superior prediction accuracy: the overall Gauge MAE is 1.417meV1.417\ $\mathrm${meV}, closely matching the ground-truth labels obtained from DFT calculations.
In Fig. 3, we report a fine-grained evaluation of prediction accuracy by partitioning the test set into subsets defined by chemical elements. For each element, we gather all crystal structures that contain it and compute the mean error within this subset. The resulting per-element statistics are visualized on the periodic table, providing a clear view of how the model generalizes across chemically diverse systems. The analysis shows that for the most of elements the prediction errors are below 1.5meV1.5\ $\mathrm${meV}, confirming the robustness of our approach across a broad spectrum of the periodic table.
In material systems, the band structure serves as a fundamental representation of the electronic structure. We sample a variety of diverse examples from the testing set and compared the band structures derived from our predicted Hamiltonians with those obtained from DFT. The results show that the band structures predicted by our method closely match those obtained from DFT, highlighting its strong potential for practical applications in materials science. Details are provided in Appendix A.6. As detailed in Appendix A.7, we evaluate the end-to-end inference time of our method, including both the construction of the zeroth-step Hamiltonian and the neural network inference, and compare it with the DFT method. The results show that our approach delivers a substantial speedup.
As detailed in Appendix A.8, we conduct fine-grained ablation studies to assess the contributions of individual components in our framework. The results demonstrate that the physics-informed input descriptor $\mathbf{H}^{(0)}$, the correction-based regression target design, the TraceGrad mechanism, the ensemble strategy, and the joint $R$- and $k$-space training objective each provide significant error reductions, and that their combination corresponds to the best overall performance. In particular, incorporating the $k$-space loss markedly enhances the accuracy of the predicted band structures and effectively suppresses unphysical artifacts such as ghost states.
Figure 3: Element-wise analysis of prediction errors. For each chemical element, we collect all of the testing structures containing that element and compute the Gauge MAE values for each subset.
6 Conclusion
We advance universal Hamiltonian deep learning through both a new method and a new dataset . We propose NextHAM, a unified deep learning framework designed for accurate and generalizable prediction of electronic-structure Hamiltonians across the periodic table. First, we leverage zeroth-step Hamiltonians, constructed from initial charge densities, as both informative input features and output priors, allowing the model to perform correction-based regression and significantly reducing the learning complexity. Second, we present a Transformer-based neural architecture that enforces strict $E(3)$-equivariance while maintaining high expressive capacity, enabling accurate modeling of spatial symmetries in material systems. Third, we design a novel training objective that jointly optimizes the Hamiltonian prediction in both real space and reciprocal space, ensuring consistency with downstream physical quantities such as band structures. We also release Materials-HAM-SOC, a diverse-collection benchmark of 17,000 DFT-calculated material structures spanning six rows of the periodic table, with explicit spin–orbit coupling and high-resolution orbital representations, providing high-quality resources for training and evaluation. Empirically, NextHAM attains DFT-level accuracy for Hamiltonians and band structures while bringing substantial speedups over conventional DFT workflows, providing powerful tools to efficient simulation and design of new materials.
Appendix A Appendix
A.1 Electronic Structure Calculations: from Density Functional Theory to Deep Learning Methods
Density Functional Theory (DFT) (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) has established itself as a foundational tool in modern electronic structure theory, with wide-ranging applications in condensed matter physics, quantum chemistry, and materials science. First developed in the 1960s by Hohenberg, Kohn, and Sham, DFT reformulates the many-electron problem by replacing the complex many-body wavefunction with the electron density $\rho(\mathbf{r})$ as the central variable. This shift dramatically simplifies the computational treatment of quantum systems while retaining the essential physics, making it feasible to study realistic systems under accepted computational cost. Over the years, DFT has become indispensable for tasks such as computing band structures and orbital energies, performing structural optimizations, and predicting a variety of electronic, magnetic, and optical properties. Its broad applicability and computational efficiency have cemented its role as a key methodology across multiple scientific domains.
At the heart of density functional theory (DFT) lies the Kohn–Sham (KS) equation (Kohn & Sham, 1965), which reformulates the many-body electronic problem into a tractable set of single-particle equations:
$\hat{H}\psi_{i}(\mathbf{r})=\epsilon_{i}\psi_{i}(\mathbf{r}),\quad\text{with}\quad\hat{H}=-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\text{ext}}(\mathbf{r})+V_{\text{HXC}}[\rho](\mathbf{r}),$ (5)
where $\hat{H}$ is the effective single-particle Hamiltonian. The potential includes the external potential $V_{\text{ext}}(\mathbf{r})$, and the Hartree–exchange–correlation (HXC) potential $V_{\text{HXC}}[\rho](\mathbf{r})=V_{\text{H}}[\rho](\mathbf{r})+V_{\text{XC}}[\rho](\mathbf{r})$, which is a functional of the electron density $\rho(\mathbf{r})$. The density itself is obtained from the KS orbitals via:
$\rho(\mathbf{r})=\sum_{m=1}^{M}|\psi_{m}(\mathbf{r})|^{2},$ (6)
where $M$ is the number of occupied single-particle states.
To numerically solve Eq. (5), a basis set is introduced. Atomic orbitals (Lin et al., 2023) are a widely adopted choice due to their localized nature and computational efficiency—they typically require fewer basis functions to reach a given level of accuracy compared to plane-wave or other delocalized bases. The atomic basis functions are products of a radial function and a spherical harmonic, that is,
[…]
where $\kappa$ denotes the element type, $lm$ denotes the angular momentum and the magnetic quantum number. Usually, real spherical harmonic functions are used. The radial functions are typically tabulated numerically on a fine radial mesh, and hence these basis functions are referred to as NAOs. the radial functions $f_{\kappa\zeta l}(\mathbf{r})$ are expanded in terms of spherical Bessel functions and truncated beyond a cutoff distance $r_{c}$
$f_{\kappa\zeta l}(\mathbf{r})=\begin{cases}\displaystyle\sum_{q}c_{\kappa\zeta lq}j_{l}(qr),&r<r_{c},\\[6.0pt] 0,&r\geq r_{c}.\end{cases}$ (8)
The KS eigenfunctions are expanded in terms of these atomic orbitals:
$\psi_{n\mathbf{k}}(\mathbf{r})=\frac{1}{\sqrt{N_{k}}}\sum_{\mathbf{R}}\sum_{\mu}C_{n\alpha,\mathbf{k}}e^{i\mathbf{k}\cdot\mathbf{R}}\phi_{u}(\mathbf{r}-\tau_{i}-\mathbf{R}),$ (9)
where $\phi_{u}(\mathbf{r}-\tau_{i}-\mathbf{R})$ are the $u$th atomic orbitals centered on the $i$th atom in the unit cell $\mathbf{R}$, and $\alpha=\{u,i\}$ is the composite index for the NAOs. $C_{n\alpha,\mathbf{k}}$ are the coefficients of orbitals $\alpha$ of band $n$ at $\mathbf{k}$ point, and $N_{k}$ is the number of unit cells in the Born–von–Kármán supercell under the periodic boundary conditions, equivalent to the number of $\mathbf{k}$ points in the first Brillouin zone (BZ).
Given the expansion of the KS states in terms of atomic orbitals in Eq. (9), the KS Eq. (5) becomes ageneralized eigenvalue problem,
$H(\mathbf{k})C_{\mathbf{k}}=E_{\mathbf{k}}S(\mathbf{k})C_{\mathbf{k}},$ (10)
where $H(\mathbf{k})$, $S(\mathbf{k})$, and $C_{\mathbf{k}}$ are the Hamiltonian matrix, overlap matrix and eigenvectors at a given $\mathbf{k}$ point, respectively. $E_{\mathbf{k}}$ is a diagonal matrix whose entries are the KS eigenenergies, $\epsilon_{n\mathbf{k}}$ denotes the energy eigenvalue of the $n$-th KS eigenstate. To obtain the Hamiltonian matrix $H(\mathbf{k})$, we first calculate the Hamiltonian in real space as
$H_{\alpha\beta}(\mathbf{R})=\left\langle\phi_{\alpha 0}\middle|-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\text{ext}}+V_{\text{HXC}}[\rho]\middle|\phi_{\beta\mathbf{R}}\right\rangle,$ (11)
where $\alpha,\beta$ are atomic orbital indices within one unit cell, and $\phi_{\alpha 0}\overset{\mathrm{def}}{=}\phi_{u}(\mathbf{r}-\tau_{i})$, $\phi_{\beta\mathbf{R}}\overset{\mathrm{def}}{=}\phi_{v}(\mathbf{r}-\tau_{j}-\mathbf{R})$. The Hamiltonian matrix at a given $\mathbf{k}$ point can be obtained via a Fourier transform,
$H_{\alpha\beta}(\mathbf{k})=\sum_{\mathbf{R}}e^{i\mathbf{k}\cdot\mathbf{R}}H_{\alpha\beta}(\mathbf{R}).$ (12)
Similarly, the overlap matrix at a given $\mathbf{k}$ point is obtained as
$S_{\alpha\beta}(\mathbf{k})=\sum_{\mathbf{R}}e^{i\mathbf{k}\cdot\mathbf{R}}S_{\alpha\beta}(\mathbf{R}),$ (13)
where
$S_{\alpha\beta}(\mathbf{R})=\langle\phi_{\alpha 0}|\phi_{\beta\mathbf{R}}\rangle.$ (14)
The overall computational procedure follows an iterative self-consistent (SC) loop:
  1. 1.
    Initial Guess: Start with an initial electron density $\rho^{(0)}(\mathbf{r})$. Initialize the number of iterations $t$ to 0.
  2. 2.
    Potential Construction: Compute the effective potential $V_{\text{HXC}}^{(t)}[\rho](\mathbf{r})$ by $\rho^{(t)}(\mathbf{r})$.
  3. 3.
    Hamiltonian Assembly: Construct the Hamiltonian matrix $\mathbf{H}^{(t)}$ using the current potential using Eq. (11).
  4. 4.
    Eigenproblem Solution: Perform a Fourier transformation and solve the generalized eigenvalue problem in Eq. (10) to obtain the KS eigenfunctions $\psi_{n\mathbf{k}}(\mathbf{r})$ and eigenvalues $\epsilon_{n\mathbf{k}}$.
  5. 5.
    Density Update: Compute the updated density $\rho^{(t)}(\mathbf{r})$ from the new orbitals using Eq. (6).
  6. 6.
    Convergence Check: Update $t\rightarrow t+1$, repeat steps 2–6 until the input and output densities agree within a chosen convergence threshold.
This procedure can be summarized schematically as:
$\rho^{(0)}(\mathbf{r})\rightarrow V^{(0)}_{\text{HXC}}[\rho](\mathbf{r})\rightarrow\mathbf{H}^{(0)}\rightarrow\psi^{(0)}_{n\mathbf{k}}(\mathbf{r})\rightarrow\rho^{(1)}(\mathbf{r})\rightarrow\cdots\rightarrow\rho^{(T)}(\mathbf{r})\rightarrow V^{(T)}_{\text{HXC}}[\rho](\mathbf{r})\rightarrow\mathbf{H}^{(T)}$
Once self-consistency is reached at iteration $T$, the final Hamiltonian matrix $\mathbf{H}^{(T)}$ can be used to compute down-stream physical quantities such as total energy, band structure, orbital energies, and derived electronic, magnetic, or transport properties.
Despite the remarkable success of Kohn–Sham DFT in advancing fields such as materials science, energy, and biomedicine over recent decades (Nagy, 1998; Jones, 2015), it still faces significant computational challenges, especially when applied to large atomic systems under limited computational resources. The primary bottlenecks arise from two aspects. First, the matrix diagonalization in Eq. (10) scales as $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in the system. Second, the iterative nature of the SC procedure requires $T$ rounds of self-consistent updates, which further amplifies the overall computational cost. This becomes particularly problematic when a high level of convergence accuracy is needed or when dealing with complex systems, often making it difficult to complete the calculations within reasonable time or resource constraints.
To address this challenge, recent approaches (Li et al., 2022; Gong et al., 2023; Yu et al., 2023; Li et al., 2025; Luo et al., 2025; Yin et al., 2025) have adopted the deep graph learning paradigm to predict the self-consistent Hamiltonians. These methods bypass the iterative and computationally intensive matrix diagonalization steps in traditional DFT algorithms by directly predicting the final converged Hamiltonian matrix $H^{(T)}_{\alpha\beta}$ in a single forward pass. As shown in Eq. (11), the Hamiltonian matrix is inherently sparse: only pairs of atoms within a cutoff radius contribute non-zero elements. Therefore, the total number of Hamiltonian matrix elements that need to be computed scales with the number of local atomic pairs in the system, leading to a complexity of $\mathcal{O}(N\overline{E})$, where $N$ is the total number of atoms and $\overline{E}$ denotes the average number of neighboring atoms within the cutoff radius per atom. Since the atomic orbital basis functions have finite spatial support, matrix elements vanish beyond a certain inter-atomic distance. In small systems where all atoms lie within each other’s cutoff radius, $\overline{E}\sim N$, and the total number of non-zero elements scales as $\mathcal{O}(N^{2})$. However, in sufficiently large systems, $\overline{E}$ saturates to a constant determined by local geometry, making the number of non-zero Hamiltonian elements scale linearly as $\mathcal{O}(N)$. Moreover, since most physical properties, such as transport, optical, and topological properties, depend only on the energy bands near the Fermi level, it is unnecessary to solve for the eigenfunctions of all occupied states once the Hamiltonian is known. Since the Hamiltonian matrix is sparse and only a limited number of bands near the Fermi level are needed, these eigenstates can be efficiently computed using methods like the shift-invert approach available in the ARPACK package (Lehoucq et al., 1998), with a computational complexity of $\mathcal{O}(N)$ for large systems.
Recently, deep-learning approaches (Li et al., 2022; Gong et al., 2023; Yu et al., 2023; Li et al., 2025; Luo et al., 2025; Yin et al., 2025) exploit the sparsity of the Hamiltonian, yielding a computational cost that scales approximately linearly with the number of non-zero matrix elements and enabling efficient, scalable prediction of quantum properties in large atomic systems. As a result, they offer a significant efficiency advantage over traditional DFT methods with computational complexity of $\mathcal{O}(TN^{3})$. This efficiency makes them particularly promising for predicting electronic structures of complex atomic systems under limited computational resources, potentially accelerating down-stream application areas like materials simulation and design.
A.2 Introduction of Reciprocal Space Electronic-Structure Hamiltonians into Deep Learning Paradigm
The self-consistent Hamiltonian $\mathbb{H}^{(T)}$, obtained through the procedure described in Section A.1, is inherently defined in real space. Its matrix elements $H_{\alpha\beta}^{(T)}$ are constructed over localized atomic orbital basis functions centered at atoms, and are truncated beyond a spatial cutoff. While real-space representations are efficient for representing local interactions, many physical phenomena such as band structures, effective low-energy models, and quasiparticle dynamics are most naturally described in reciprocal space.
To obtain a reciprocal-space Hamiltonian, we perform a Fourier transformation of the real-space matrix elements. For a periodic system with lattice vectors $\{\mathbf{R}\}$, the Bloch Hamiltonian $H(\mathbf{k})$ at wavevector $\mathbf{k}$ is defined as:
$H_{\alpha\beta}(\mathbf{k})=\sum_{\mathbf{R}}e^{i\mathbf{k}\cdot\mathbf{R}}H_{\alpha\beta}(\mathbf{R}),$ (15)
where $i$ is the imaginary unit ($i^{2}=-1$), and $H_{\alpha\beta}(\mathbf{R})$ denotes the real-space Hamiltonian matrix element between orbital $\alpha$ in a reference unit cell and orbital $\beta$ in a cell displaced by lattice vector $\mathbf{R}$. These elements are directly taken from the converged real-space Hamiltonian $\mathbf{H}^{(T)}$ defined over the localized atomic orbital basis, with each pair of orbitals uniquely associated with a displacement vector $\mathbf{R}$. For simplicity, we omit the superscript $(T)$ in Eq. (15), with the understanding that all real-space matrix elements originate from $\mathbf{H}^{(T)}$.
Diagonalizing $H(\mathbf{k})$ at each wavevector $\mathbf{k}$ in the Brillouin zone yields the system’s electronic band structure:
$H(\mathbf{k})\psi_{n\mathbf{k}}=\varepsilon_{n\mathbf{k}}S(\mathbf{k})\psi_{n\mathbf{k}},$ (16)
Let $\hat{H}(\mathbf{k})\in\mathbb{C}^{n\times n}$ denote a Hermitian matrix that approximates $H(\mathbf{k})$. It can also be solved through a generalized eigenvalue equation to obtain the eigenvalues and wave functions.
$\hat{H}(\mathbf{k})\hat{\psi}_{n\mathbf{k}}=\hat{\varepsilon}_{n\mathbf{k}}S(\mathbf{k})\hat{\psi}_{n\mathbf{k}},$ (17)
where $\varepsilon_{n\mathbf{k}}$ and $\hat{\varepsilon}_{n\mathbf{k}}$ are diagonal matrices of eigenvalues, and $\psi_{n\mathbf{k}}$ and $\hat{\psi}_{n\mathbf{k}}$ are the corresponding eigenvectors. Define $\Delta H(\mathbf{k})=H(\mathbf{k})-\hat{H}(\mathbf{k})$. Assume a spectral gap $\delta$ separates the generalized eigenvalues of $H(\mathbf{k})$ and $\hat{H}(\mathbf{k})$. $\kappa(\cdot)$ denotes the condition number of a given matrix, $\|\cdot\|_{2}$ represents the spectral norm, $\|\Delta H(\mathbf{k})\|_{1,1}=\sum_{i,j}|\Delta H_{ij}(\mathbf{k})|$. Then, the difference in eigenvalues and the angle $\theta$ between the eigenspace of $H(\mathbf{k})$ and $\hat{H}(\mathbf{k})$ satisfy:
  1. 1.
    Eigenvalue Differences:
    $\bigl|\varepsilon_{n\mathbf{k}}-\hat{\varepsilon}_{n\mathbf{k}}|\leq\frac{\kappa(S(\mathbf{k}))}{\|S(\mathbf{k})\|_{2}}\,\|\Delta H(\mathbf{k})\|_{1,1},$
  2. 2.
    Eigenspace Angle:
    $\sin\theta\leq\frac{\kappa(S(\mathbf{k}))}{\|S(\mathbf{k})\|_{2}}\,\frac{\|\Delta H(\mathbf{k})\|_{1,1}}{\delta},$
where $\theta$ is the angle between the eigenspaces corresponding to $\varepsilon_{n\mathbf{k}}$ and $\hat{\varepsilon}_{n\mathbf{k}}$. The theorem(Golub & Van Loan, 2013) highlights that due to the non-orthogonality of the orbital basis set, the errors in band energies and wave functions can be amplified by the condition number factor $\frac{\kappa(S(\mathbf{k}))}{\|S(\mathbf{k})\|_{2}}$. As a result, even a small error may cause the band eigenvalues and wave functions to deviate significantly from the true results, manifesting as the appearance of ghost states in the band structure.
To mitigate the amplification of perturbations in the predicted results caused by the condition number, a feasible approach is to perform a basis transformation for the Hamiltonian matrix $H(\mathbf{k})$ by introducing projection operators $\mathcal{U}(\mathbf{k})$ formed from the complete set of eigenstates $\psi_{n\mathbf{k}}$}, thereby transforming $H(\mathbf{k})$ into a diagonal representation. From a physical perspective, the low-energy subspace near the Fermi level governs essential material properties such as optical, thermal and transport behaviors. Accordingly, the projected Hamiltonian $H(\mathbf{k})$ can be decomposed into three parts of the projection space:
  • Low-energy subspace $\widetilde{\mathbf{H}}_{PP}(\mathbf{k})$: $\mathcal{P}(\mathbf{k})$ are spanned by $N_{P}$ eigenvectors $\{\psi_{n\mathbf{k}}\}$ with energies below the cutoff energy, the $H(\mathbf{k})$ are projected into $\mathcal{P}(\mathbf{k})$ space.
  • High-energy subspace $\widetilde{\mathbf{H}}_{QQ}(\mathbf{k})$: $\mathcal{Q}(\mathbf{k})$ are spanned by the remaining $N_{Q}$ eigenvectors above the cutoff energy, the $H(\mathbf{k})$ are projected into $\mathcal{Q}(\mathbf{k})$ space.
  • Coupling subspace $\widetilde{\mathbf{H}}_{PQ}(\mathbf{k})$: the off-diagonal coupling between $P$ and $Q$, encoded in the cross blocks of the full Hamiltonian.
Let $\mathcal{P}(\mathbf{k})\in\mathbb{C}^{N\times N_{P}}$ and $\mathcal{Q}(\mathbf{k})\in\mathbb{C}^{N\times N_{Q}}$ be the matrices whose columns are orthonormal eigenvectors spanning the low- and high-energy subspaces, stacking the bases as
$\mathcal{U}(\mathbf{k})=\big[\,\mathcal{P}(\mathbf{k})\;\mathcal{Q}(\mathbf{k})\,\big]\in\mathbb{C}^{N\times(N_{P}+N_{Q})},$
and assuming $N_{P}+N_{Q}=N$ ($\mathcal{U}(\mathbf{k})^{\dagger}S(\mathbf{k})\mathcal{U}(\mathbf{k})=\mathbf{1}$), the Hamiltonian in the $(\mathcal{P},\mathcal{Q})$ basis is obtained by a single similarity transform:
$\widetilde{\mathbf{H}}(\mathbf{k})\;=\;\mathcal{U}(\mathbf{k})^{\dagger}\,\mathbf{H}(\mathbf{k})\,\mathcal{U}(\mathbf{k})\;=\;\begin{bmatrix}\widetilde{\mathbf{H}}_{PP}(\mathbf{k})&\widetilde{\mathbf{H}}_{PQ}(\mathbf{k})\\ \widetilde{\mathbf{H}}_{QP}(\mathbf{k})&\widetilde{\mathbf{H}}_{QQ}(\mathbf{k})\end{bmatrix}$.
For the ground-truth Hamiltonian, when transformed by its own eigenbasis $\mathcal{U}(\mathbf{k})$, the cross block vanishes, i.e., $\widetilde{\mathbf{H}}_{PQ}(\mathbf{k})=\mathbf{0}$. In contrast, when a predicted Hamiltonian is projected onto the subspaces defined by the ground-truth eigenbasis, the mismatch between the predicted and exact eigenvectors may produce spurious non-zero entries, $\widetilde{\mathbf{H}}_{PQ}(\mathbf{k})\neq 0$. These unphysical couplings manifest as artifacts such as ghost states, and thus provide a meaningful signal for penalization during training.
Because the eigenvalues of $\mathbf{H}(\mathbf{k})$ directly define the band structure, reciprocal-space supervision provides a natural training signal. We therefore assign distinct loss terms to the three components. The low-energy block $\widetilde{\mathbf{H}}_{PP}(\mathbf{k})$ governs the states near the Fermi level and thus dominates observable physics; accurate supervision on this block is crucial. The high-energy block $\widetilde{\mathbf{H}}_{QQ}(\mathbf{k})$ does not directly determine low-energy phenomena, but maintaining its fidelity is important: otherwise errors in $Q$ may propagate indirectly through erroneous $PQ$ couplings. Finally, the cross block $\widetilde{\mathbf{H}}_{PQ}(\mathbf{k})$ should ideally vanish; we enforce this by adding an explicit penalty on $\lVert\widetilde{\mathbf{H}}_{PQ}(\mathbf{k})\rVert$, which suppresses unphysical couplings between $P$ and $Q$, thereby eliminating ghost states and restoring the intended decoupling of subspaces.
A.3 Details on Training Loss Functions
We elaborate on the details of Eq. (1) in the following equation:
$\begin{split}&\mathrm{loss}(\mathbf{R})={\mathbb{E}}_{\mathbf{R}}[\lambda_{R}\Big((1-\lambda_{C})\cdot\mathrm{loss}_{H}(\mathbf{R})+\gamma(\mathrm{loss}_{H},\mathrm{loss}_{T},\lambda_{C})\cdot\mathrm{loss}_{T}(\mathbf{R})\Big)],\\ &\mathrm{loss}_{H}(\mathbf{R})=MSE\big(\widehat{\mathbf{H}}(\mathbf{R}),\mathbf{H}^{gt}(\mathbf{R},\mu)\big),\\ &\mathrm{loss}_{T}(\mathbf{R})=MAE\big(\widehat{\mathbf{T}}(\mathbf{R}),\mathbf{T}^{gt}(\mathbf{R},\mu)\big),\\ &\gamma(\mathrm{loss}_{H},\mathrm{loss}_{T},\lambda_{C})=\lambda_{C}\cdot\mathrm{No\_Grad}\left(\frac{\mathrm{loss}_{H}(\mathbf{R})}{\mathrm{loss}_{T}(\mathbf{R})}\right).\end{split}$ (18)
where $\lambda_{R}$ is a hyper-parameter, $\mathbf{R}$ denotes the lattice vector connecting the reference unit cell and a neighboring unit cell. $\widehat{\mathbf{H}}(\mathbf{R})$ and $\widehat{\mathbf{T}}(\mathbf{R})$ denote the predicted Hamiltonian and its corresponding trace quantity in real space, respectively. Here, we compute $\widehat{\mathbf{H}}(\mathbf{R})$ as:
$\widehat{\mathbf{H}}(\mathbf{R})=\mathbf{H}^{(0)}(\mathbf{R})+\widehat{\Delta\mathbf{H}}(\mathbf{R}),$
where $\widehat{\Delta\mathbf{H}}(\mathbf{R})$ is the predicted correction term of the Hamiltonian.
The ground truth Hamiltonians are denoted as $\mathbf{H}^{gt}(\mathbf{R})=\mathbf{H}^{(T)}(\mathbf{R})$. However, rather than directly using these ground truth values to supervise $\widehat{\mathbf{H}}(\mathbf{R})$, we construct augmented supervision targets by introducing an additional term:
$\mathbf{H}^{gt}(\mathbf{R},\mu)=\mathbf{H}^{gt}(\mathbf{R})+\mu\cdot\mathbf{S}(\mathbf{R}),$ (19)
where $\mu$ is a scalar coefficient, $\mathbf{S}(\mathbf{R})$ denotes the real-space overlap matrix, and $\mathbf{H}^{(0)}(\mathbf{R})$ denotes the real-space zeroth-step Hamiltonian matrix. Following the gauge-error formulation of Wang et al. (2024b), adding a shift term $\mu\cdot\mathbf{S}(\mathbf{R})$ to the Hamiltonian leaves all down-stream physical observables unchanged. In practice, $\mu$ is chosen as the solution that minimizes the overall loss, as established in Wang et al. (2024b). This removes the gauge freedom inherent in the Hamiltonian representation, facilitating more stable and efficient convergence of the neural network.
The corresponding trace quantity used as the supervision signal is computed as:
$\mathbf{T}^{gt}(\mathbf{R},\mu)$ $=\mathrm{tr}\left(\Delta\mathbf{H}^{gt}(\mathbf{R},\mu)\cdot\Delta\mathbf{H}^{gt}(\mathbf{R},\mu)^{\dagger}\right)$ (20)
$=\mathrm{tr}\left((\mathbf{H}^{gt}(\mathbf{R},\mu)-\mathbf{H}^{(0)}(\mathbf{R}))\cdot(\Delta\mathbf{H}^{gt}(\mathbf{R},\mu)-\mathbf{H}^{(0)}(\mathbf{R}))^{\dagger}\right),$
where $\mathbf{H}^{(0)}(\mathbf{R})$ denotes the real-space zeroth-step Hamiltonian matrix.
Inspired by Yin et al. (2025), the scaling factor $\gamma(\mathrm{loss}_{H},\mathrm{loss}_{T},\lambda_{C})$ in Eq. (LABEL:loss_real_detail) is designed to harmonize the contributions from the two loss terms, ensuring stable optimization. Here, $\lambda_{C}$ is a hyper-parameter that controls the overall strength of the balancing mechanism. The term $\mathrm{No\_Grad}(\cdot)$ ensures that gradients are dropped during the computation of this coefficient, preventing interference with the back-propagation of $\mathrm{loss}_{T(R)}$. By applying this balancing strategy,better numerical stability and balanced learning performance across both the Hamiltonian and trace quantity supervision branches can be achieved.
We elaborate on the details of Eq. (2) as follows. Let $\widehat{\mathbf{H}}(\mathbf{k})$ denote the predicted full Hamiltonian in reciprocal space, obtained from the Fourier transform of $\widehat{\mathbf{H}}(\mathbf{R})$ using Eq. (15). Similarly, let $\mathbf{H}^{gt}(\mathbf{k},\mu)$ denote the ground-truth Hamiltonian in reciprocal space, obtained from the Fourier transform of $\mathbf{H}^{gt}(\mathbf{R},\mu)$. Both Hamiltonians are projected by the ground-truth eigenbasis $\mathcal{U}(\mathbf{k})=[\,\mathcal{P}(\mathbf{k}),\,\mathcal{Q}(\mathbf{k})\,]$, yielding block-partitioned forms:
$\widetilde{\mathbf{H}}(\mathbf{k})=\mathcal{U}(\mathbf{k})^{\dagger}\,\widehat{\mathbf{H}}(\mathbf{k})\,\mathcal{U}(\mathbf{k})=\begin{bmatrix}\widetilde{\mathbf{H}}_{PP}(\mathbf{k})&\widetilde{\mathbf{H}}_{PQ}(\mathbf{k})\\ \widetilde{\mathbf{H}}_{QP}(\mathbf{k})&\widetilde{\mathbf{H}}_{QQ}(\mathbf{k})\end{bmatrix}$,
$\widetilde{\mathbf{H}}^{gt}(\mathbf{k},\mu)=\mathcal{U}(\mathbf{k})^{\dagger}\,\mathbf{H}^{gt}(\mathbf{k},\mu)\,\mathcal{U}(\mathbf{k})=\begin{bmatrix}\widetilde{\mathbf{H}}_{PP}^{gt}(\mathbf{k},\mu)&\widetilde{\mathbf{H}}_{PQ}^{gt}(\mathbf{k},\mu)\\ \widetilde{\mathbf{H}}_{QP}^{gt}(\mathbf{k},\mu)&\widetilde{\mathbf{H}}_{QQ}^{gt}(\mathbf{k},\mu)\end{bmatrix}$.
For the exact Hamiltonian, the off-diagonal block ideally vanishes, i.e., $\widetilde{\mathbf{H}}_{PQ}^{gt}(\mathbf{k},\mu)=\mathbf{0}$, whereas for the predicted Hamiltonian, spurious non-zero entries generally appear in $\widetilde{\mathbf{H}}_{PQ}(\mathbf{k})$, manifesting as unphysical ghost states.
The loss is then defined block-wise:
$\begin{split}\mathrm{loss}(\mathbf{k})&=\mathbb{E}_{\mathbf{k}}\Big[\lambda_{\mathcal{P}}\cdot\mathrm{MSE}\big(\widetilde{\mathbf{H}}_{PP}(\mathbf{k}),\widetilde{\mathbf{H}}_{PP}^{gt}(\mathbf{k},\mu)\big)\\ &\quad+\lambda_{\mathcal{Q}}\cdot\mathrm{MSE}\big(\widetilde{\mathbf{H}}_{QQ}(\mathbf{k}),\widetilde{\mathbf{H}}_{QQ}^{gt}(\mathbf{k},\mu)\big)\\ &\quad+\lambda_{PQ}\cdot\mathrm{MSE}\big(\widetilde{\mathbf{H}}_{PQ}(\mathbf{k}),\widetilde{\mathbf{H}}_{PQ}^{gt}(\mathbf{k},\mu)\big)\Big].\end{split}$ (21)
where $\lambda_{\mathcal{P}},\lambda_{\mathcal{Q}},\lambda_{PQ}$ are tunable hyperparameters controlling the relative importance of the three terms.
The overall loss function combines the losses from both R-space and k-space:
$\begin{split}\mathrm{loss}_{all}=&\mathrm{loss}(\mathbf{R})+\mathrm{loss}(\mathbf{k})\\ =&{\mathbb{E}}_{\mathbf{R}}[\lambda_{R}\Big((1-\lambda_{C})\cdot\mathrm{loss}_{H}(\mathbf{R})+\gamma(\mathrm{loss}_{H},\mathrm{loss}_{T},\lambda_{C})\cdot\mathrm{loss}_{T}(\mathbf{R})\Big)]\\ &+{\mathbb{E}}_{\mathbf{k}}[\lambda_{\mathcal{P}}\cdot\mathrm{loss}_{\mathcal{P}}(\mathbf{k})+\lambda_{\mathcal{Q}}\cdot\mathrm{loss}_{\mathcal{Q}}(\mathbf{k})+\lambda_{PQ}\cdot\mathrm{loss}_{PQ}(\mathbf{k})]\end{split}$ (22)
where the value of $\mu$ is determined by $\frac{\partial\mathrm{loss}_{all}}{\partial\mu}=0$. It can be solved analytically as:
[…]
which obtains:
$\mu=\frac{\Delta_{1}}{\Delta_{2}},$ (24)
$\Delta_{1}=\frac{\lambda_{R}}{\text{N}_{R}}\sum_{\mathbf{R},\alpha\beta}\text{Re}\Bigl([\widehat{\mathbf{H}}(\mathbf{R})-\mathbf{H}^{gt}(\mathbf{R})]^{*}_{\alpha\beta}\mathbf{S}(\mathbf{R})_{\alpha\beta}\Bigr)$
$\,\,\,\,\,\,\,\,\,+\frac{\lambda_{\mathcal{P}}}{\text{N}_{\mathcal{P}}}\sum_{\mathbf{k},\alpha}[\widetilde{\mathbf{H}}_{PP}(\mathbf{k})-\widetilde{\mathbf{H}}^{gt}_{PP}(\mathbf{k})]_{\alpha\alpha}$
$\,\,\,\,\,\,\,\,\,+\frac{\lambda_{\mathcal{Q}}}{\text{N}_{\mathcal{Q}}}\sum_{\mathbf{k},\alpha}[\widetilde{\mathbf{H}}_{QQ}(\mathbf{k})-\widetilde{\mathbf{H}}^{gt}_{QQ}(\mathbf{k})]_{\alpha\alpha}$
$\Delta_{2}=\frac{\lambda_{R}}{\text{N}_{R}}\sum_{\mathbf{R},\alpha\beta}\mathbf{S}(\mathbf{R})^{*}_{\alpha\beta}\mathbf{S}(\mathbf{R})_{\alpha\beta}+\sum_{\mathbf{k},\alpha}\frac{\lambda_{\mathcal{P}}}{\text{N}_{\mathcal{P}}}+\sum_{\mathbf{k},\alpha}\frac{\lambda_{\mathcal{Q}}}{\text{N}_{\mathcal{Q}}}$
where * denotes the complex conjugate operation, $\text{N}_{R}$, $\text{N}_{\mathcal{P}}$, $\text{N}_{\mathcal{Q}}$, and $\text{N}_{\mathcal{PQ}}$ denote the total number of Hamiltonian matrix elements corresponding to real space, $\mathcal{P}$ space, $\mathcal{Q}$ space, and their coupling space respectively. $\widetilde{\mathbf{H}}^{gt}_{PP}(\mathbf{k})$ and $\widetilde{\mathbf{H}}^{gt}_{QQ}(\mathbf{k})$ are computed from the ground-truth Hamiltonian $\mathbf{H}^{gt}(\mathbf{k})$ by:
$\mathcal{U}(\mathbf{k})^{\dagger}\,\mathbf{H}^{gt}(\mathbf{k})\,\mathcal{U}(\mathbf{k})=\begin{bmatrix}\widetilde{\mathbf{H}}^{gt}_{PP}(\mathbf{k})&\mathbf{0}\\ \mathbf{0}&\widetilde{\mathbf{H}}^{gt}_{QQ}(\mathbf{k})\end{bmatrix}$.
It is important to clarify that, in the analytical derivation of $\mu$ in Eq.(23), the real-space contribution can be written in simplified form as:
$\lambda_{R}\cdot\mathrm{loss}_{H}(\mathbf{R})=\frac{\lambda_{R}}{\text{N}_{R}}\sum_{\mathbf{R},\alpha\beta}\Big[\Big|\Big(\widehat{\mathbf{H}}(\mathbf{R})-\mathbf{H}^{gt}(\mathbf{R})\Big)_{\alpha\beta}\Big|^{2}+\mu^{2}\Big|\mathbf{S}(\mathbf{R})_{\alpha\beta}\Big|^{2}-2\mu\text{Re}\Bigl([\widehat{\mathbf{H}}(\mathbf{R})-\mathbf{H}^{gt}(\mathbf{R})]^{*}_{\alpha\beta}\mathbf{S}(\mathbf{R})_{\alpha\beta}\Bigr)\Big]$
rather than explicitly retaining the trace supervision term $\mathrm{loss}_{T}(\mathbf{R})$ like:
$\lambda_{R}\cdot\Big((1-\lambda_{C})\cdot\mathrm{loss}_{H}(\mathbf{R})+\gamma(\mathrm{loss}_{H},\mathrm{loss}_{T},\lambda_{C})\cdot\mathrm{loss}_{T}(\mathbf{R})\Big).$
This simplification is purely at the algebraic and notational level and does not imply that $\mathrm{loss}_{T}(\mathbf{R})$ is omitted. In fact, the balancing factor $\gamma(\mathrm{loss}_{H},\mathrm{loss}_{T},\lambda_{C})=\lambda_{C}\cdot\mathrm{No\_Grad}\left(\frac{\mathrm{loss}_{H}(\mathbf{R})}{\mathrm{loss}_{T}(\mathbf{R})}\right)$ guarantees that this weighted combination of $\mathrm{loss}_{H}(\mathbf{R})$ and $\mathrm{loss}_{T}(\mathbf{R})$ is numerically equivalent to $\mathrm{loss}_{H}(\mathbf{R})$, in which a fixed fraction of the contribution has been substituted by $\mathrm{loss}_{T}(\mathbf{R})$ in a stable and adaptive manner. In other words, $\mathrm{loss}_{T}$ serves as a surrogate for a controlled fraction of $\mathrm{loss}_{H}$, while after normalization the effective value of the entire term remains consistent with $\mathrm{loss}_{H}(\mathbf{R})$. Therefore, in the derivation of $\mu$, it is sufficient and mathematically consistent to retain only $\mathrm{loss}_{H}(\mathbf{R})$, while the beneficial regularization effect of $\mathrm{loss}_{T}$ is still fully incorporated through the design of $\gamma(\cdot)$.
A.4 Dataset Details
To construct Materials-HAM-SOC, the first-principles calculations are performed using the Atomic-Orbital Based Ab-initio Computation at USTC (ABACUS)(Li et al., 2016; Lin et al., 2023) package. The Perdew–Burke–Ernzerhof (PBE) exchange-correlation functional (Perdew et al., 1996) and the optimized norm-conserving Vanderbilt (ONCV) fully relativistic pseudopotentials (Hamann, 2013) from the PseudoDojo library (van Setten et al., 2018) are used. Table LABEL:pse&pot summarizes the valence electron configurations used in the pseudopotentials and the corresponding numerical atomic orbital (NAO) basis for each element. In self-consistent calculations, the energy cutoff for wave functions is set to 120 Ry and the charge density was converged to a threshold of 1×10-6. The $\Gamma$-centered Monkhorst-Pack $6\times 6\times 6$ k-point mesh is employed for self-consistent calculations.
The crystal structures were obtained from the Materials Project database, from which a total of approximately 17,000 nonmagnetic compounds were randomly selected. Among them, 12,000 structures were used for training, 2,000 for validation, and 3,000 for testing. The statistical distributions of atomic species and atomic counts in the training, validation, and test sets are illustrated in Figures 4 and 5. Furthermore, the occurrence frequencies of different elements across the three subsets are presented in Figures 6(a)6(c).
We visualize representative crystal structures in Fig. 7, highlighting the diversity and broad coverage of our curated dataset Materials-HAM-SOC. These samples span a wide range of chemistries, crystal symmetries, and atomic complexities, illustrating the richness of the dataset and its suitability for training universal Hamiltonian prediction models.
Figure 4: Bar charts of elemental species distributions in the training, validation, and test sets.
Figure 5: Bar charts of atomic count distributions in the training, validation, and test sets.
(a) Training set
(b) Validation set
(c) Testing set
Figure 6: Statistical charts of element occurrence frequencies in the training, validation, and test sets.
Table 2: The valence electron configurations for pseudopotentials and corresponding NAOs of the elements used in this study.
Element Number Element Name Valence Electrons NAOs Cutoff Radius
001 H $1\text{s}^{1}$ 2s1p 7 a.u.
002 He $1\text{s}^{2}$ 2s1p 7 a.u.
003 Li $1\text{s}^{2}2\text{s}^{1}$ 4s1p 7 a.u.
004 Be $1\text{s}^{2}2\text{s}^{2}$ 4s1p 7 a.u.
005 B $2\text{s}^{2}2\text{p}^{1}$ 2s2p1d 7 a.u.
006 C $2\text{s}^{2}2\text{p}^{2}$ 2s2p1d 7 a.u.
007 N $2\text{s}^{2}2\text{p}^{3}$ 2s2p1d 7 a.u.
008 O $2\text{s}^{2}2\text{p}^{4}$ 2s2p1d 7 a.u.
009 F $2\text{s}^{2}2\text{p}^{5}$ 2s2p1d 7 a.u.
010 Ne $2\text{s}^{2}2\text{p}^{6}$ 2s2p1d 7 a.u.
011 Na $2\text{s}^{2}2\text{p}^{6}3\text{s}^{1}$ 4s2p1d 7 a.u.
012 Mg $2\text{s}^{2}2\text{p}^{6}3\text{s}^{2}$ 4s2p1d 7 a.u.
013 Al $3\text{s}^{2}3\text{p}^{1}$ 2s2p1d 7 a.u.
014 Si $3\text{s}^{2}3\text{p}^{2}$ 2s2p1d 7 a.u.
015 P $3\text{s}^{2}3\text{p}^{3}$ 2s2p1d 7 a.u.
016 S $3\text{s}^{2}3\text{p}^{4}$ 2s2p1d 7 a.u.
017 Cl $3\text{s}^{2}3\text{p}^{5}$ 2s2p1d 7 a.u.
018 Ar $3\text{s}^{2}3\text{p}^{6}$ 2s2p1d 7 a.u.
019 K $3\text{s}^{2}3\text{p}^{6}4\text{s}^{1}$ 4s2p1d 7 a.u.
020 Ca $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}$ 4s2p1d 7 a.u.
021 Sc $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}3\text{d}^{1}$ 4s2p2d1f 7 a.u.
022 Ti $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}3\text{d}^{2}$ 4s2p2d1f 7 a.u.
023 V $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}3\text{d}^{3}$ 4s2p2d1f 7 a.u.
024 Cr $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}3\text{d}^{4}$ 4s2p2d1f 7 a.u.
025 Mn $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}3\text{d}^{5}$ 4s2p2d1f 7 a.u.
026 Fe $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}3\text{d}^{6}$ 4s2p2d1f 7 a.u.
027 Co $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}3\text{d}^{7}$ 4s2p2d1f 7 a.u.
028 Ni $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}3\text{d}^{8}$ 4s2p2d1f 7 a.u.
029 Cu $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}3\text{d}^{9}$ 4s2p2d1f 7 a.u.
030 Zn $3\text{s}^{2}3\text{p}^{6}4\text{s}^{2}3\text{d}^{10}$ 4s2p2d1f 7 a.u.
031 Ga $3\text{d}^{10}4\text{s}^{2}4\text{p}^{1}$ 2s2p2d1f 8 a.u.
032 Ge $3\text{d}^{10}4\text{s}^{2}4\text{p}^{2}$ 2s2p2d1f 8 a.u.
033 As $4\text{s}^{2}4\text{p}^{3}$ 2s2p1d 7 a.u.
034 Se $4\text{s}^{2}4\text{p}^{4}$ 2s2p1d 7 a.u.
035 Br $4\text{s}^{2}4\text{p}^{5}$ 2s2p1d 8 a.u.
036 Kr $4\text{s}^{2}4\text{p}^{6}$ 2s2p1d 8 a.u.
037 Rb $4\text{s}^{2}4\text{p}^{6}5\text{s}^{1}$ 4s2p1d 9 a.u.
038 Sr $4\text{s}^{2}4\text{p}^{6}5\text{s}^{2}$ 4s2p1d 8 a.u.
039 Y $4\text{s}^{2}4\text{p}^{6}5\text{s}^{2}4\text{d}^{1}$ 4s2p2d1f 8 a.u.
040 Zr $4\text{s}^{2}4\text{p}^{6}5\text{s}^{2}4\text{d}^{2}$ 4s2p2d1f 7 a.u.
041 Nb $4\text{s}^{2}4\text{p}^{6}5\text{s}^{2}4\text{d}^{3}$ 4s2p2d1f 7 a.u.
042 Mo $4\text{s}^{2}4\text{p}^{6}5\text{s}^{2}4\text{d}^{4}$ 4s2p2d1f 7 a.u.
043 Tc $4\text{s}^{2}4\text{p}^{6}5\text{s}^{2}4\text{d}^{5}$ 4s2p2d1f 7 a.u.
044 Ru $4\text{s}^{2}4\text{p}^{6}5\text{s}^{2}4\text{d}^{6}$ 4s2p2d1f 7 a.u.
045 Rh $4\text{s}^{2}4\text{p}^{6}5\text{s}^{2}4\text{d}^{7}$ 4s2p2d1f 7 a.u.
046 Pd $4\text{s}^{2}4\text{p}^{6}4\text{d}^{10}$ 2s2p2d1f 7 a.u.
047 Ag $4\text{s}^{2}4\text{p}^{6}5\text{s}^{2}4\text{d}^{9}$ 4s2p2d1f 7 a.u.
048 Cd $4\text{s}^{2}4\text{p}^{6}5\text{s}^{2}4\text{d}^{10}$ 4s2p2d1f 7 a.u.
049 In $4\text{d}^{10}5\text{s}^{2}5\text{p}^{1}$ 2s2p2d1f 7 a.u.
050 Sn $4\text{d}^{10}5\text{s}^{2}5\text{p}^{2}$ 2s2p2d1f 7 a.u.
051 Sb $4\text{d}^{10}5\text{s}^{2}5\text{p}^{3}$ 2s2p2d1f 7 a.u.
052 Te $4\text{d}^{10}5\text{s}^{2}5\text{p}^{4}$ 2s2p2d1f 7 a.u.
053 I $5\text{s}^{2}5\text{p}^{5}$ 2s2p1d 7 a.u.
054 Xe $5\text{s}^{2}5\text{p}^{6}$ 2s2p1d 7 a.u.
055 Cs $5\text{s}^{2}5\text{p}^{6}6\text{s}^{1}$ 4s2p1d 8 a.u.
056 Ba $5\text{s}^{2}5\text{p}^{6}5\text{d}^{1}6\text{s}^{1}$ 4s2p2d1f 8 a.u.
072 Hf $5\text{s}^{2}5\text{p}^{6}6\text{s}^{2}5\text{d}^{2}$ 4s2p2d2f 7 a.u.
073 Ta $5\text{s}^{2}5\text{p}^{6}6\text{s}^{2}5\text{d}^{3}$ 4s2p2d2f 7 a.u.
074 W $5\text{s}^{2}5\text{p}^{6}6\text{s}^{2}5\text{d}^{4}$ 4s2p2d2f 7 a.u.
075 Re $5\text{s}^{2}5\text{p}^{6}6\text{s}^{2}5\text{d}^{5}$ 4s2p2d1f 7 a.u.
076 Os $5\text{s}^{2}5\text{p}^{6}6\text{s}^{2}5\text{d}^{6}$ 4s2p2d1f 7 a.u.
077 Ir $5\text{s}^{2}5\text{p}^{6}6\text{s}^{2}5\text{d}^{7}$ 4s2p2d1f 7 a.u.
078 Pt $5\text{s}^{2}5\text{p}^{6}6\text{s}^{2}5\text{d}^{8}$ 4s2p2d1f 7 a.u.
079 Au $5\text{s}^{2}5\text{p}^{6}6\text{s}^{2}5\text{d}^{9}$ 4s2p2d1f 7 a.u.
080 Hg $5\text{s}^{2}5\text{p}^{6}6\text{s}^{2}5\text{d}^{10}$ 4s2p2d1f 7 a.u.
081 Tl $5\text{d}^{10}6\text{s}^{2}6\text{p}^{1}$ 2s2p2d1f 7 a.u.
082 Pb $5\text{d}^{10}6\text{s}^{2}6\text{p}^{2}$ 2s2p2d1f 7 a.u.
083 Bi $5\text{d}^{10}6\text{s}^{2}6\text{p}^{3}$ 2s2p2d1f 7 a.u.
Figure 7: Representative crystal structures sampled from the Materials-HAM-SOC dataset. The examples cover diverse chemical compositions, structural patterns, and atomic configurations, demonstrating the dataset’s broad coverage across the periodic table. Such diversity ensures that the benchmark provides a comprehensive foundation for training and evaluating universal Hamiltonian prediction models.
A.5 Implementation Details
Our NextHAM framework is implemented based on PyTorch 2.2.0, E3NN 0.5.6, and CUDA 12.1. The training was conducted on a GPU cluster equipped with NVIDIA A800 GPUs, each with 80 GiB memory.
For the input of the neural network, we adopt a cutoff radius of $8.0\,\text{\AA }$ to define the neighboring range in the atomic graph. The angular relations between atoms are represented using spherical harmonics with degrees $0\leq l\leq 5$, while the interatomic distances are encoded through a Gaussian basis expansion (Gong et al., 2023) with a preset base number of 64. The Transformer network consists of four stacked basic blocks. Each block contains an E(3)-symmetry layer normalization module, an E(3)-symmetry feed-forward module, an E(3)-symmetry multi-head graph attention module, and a TraceGrad module. The internal node features $\mathbf{f}^{\text{(node)}}_{a}$ and edge features $\mathbf{f}^{\text{(edge)}}_{ab}$, $\mathbf{f}^{\prime\text{(edge)}}_{ab}$, as well as $\mathbf{o}^{\text{(edge)}}_{ab}$, are represented in direct-sum state form with the following structure:
$32\times 0e+16\times 1e+16\times 1o+8\times 2e+8\times 2o+8\times 3e+8\times 3o+8\times 4e+8\times 4o+1\times 5o+1\times 6e,$
where the coefficients preceding $\times$ denote multiplicities, and the subsequent terms indicate tensor degrees together with their corresponding even (e) or odd (o) parities. For the TraceGrad module, the constructed SO(3)-invariant feature $z^{\text{(edge)}}_{ab}$ has a dimension of $256$. On the output side, to map the network outputs from the direct-sum E(3)-symmetric tensors into Hamiltonian matrices, we employ the conversion modules provided by Gong et al. (2023), thereby ensuring the exact symmetry of the predicted results with SU(2) symmetry. We employ an ensemble of four sub-models to predict the electronic-structure Hamiltonian. The first sub-model is responsible for predicting the Hamiltonian submatrices formed by atomic pairs with interatomic distances in the range $0,1.0\,\text{Å})$, where the case of distance equal to zero corresponds to the on-site Hamiltonian (i.e., the Hamiltonian formed by an atom with itself). The second sub-model handles atomic pairs with distances in the range $1.0\,\text{Å},2.0\,\text{Å})$, the third sub-model covers the range $2.0\,\text{Å},4.0\,\text{Å})$, and the fourth sub-model addresses the range $4.0\,\text{Å},6.0\,\text{Å})$. For atomic pairs with distances greater than 6.0 Å, we found that their self-consistent Hamiltonian is almost identical to the zeroth-step Hamiltonian numerically. Therefore, for these distant atoms, we bypass the neural network correction step and use the zeroth-step Hamiltonian as the final result.
In the training stage, each card is assigned to one of the sub-models. In our training strategy, electronic states $\leq 10$ eV above the Fermi level are included in the low-energy subspace $\mathcal{P}$, while the remaining states are divided to the high-energy subspace $\mathcal{Q}$. We train the model for a total of $100$ epochs on the training set and evaluate the checkpoint that achieves the best performance on the validation set for testing. The hyper-parameters for loss functions are set as $\lambda_{C}=0.2$, $\lambda_{R}=0.99955$, $\lambda_{\mathcal{P}}=0.0002$, $\lambda_{\mathcal{Q}}=0.0001$, and $\lambda_{PQ}=0.00015$, determined according to the performance on the validation set. We adopt the Adam optimizer with an initial learning rate of $5\times 10^{-4}$. A warm-up phase of 5 epochs linearly increases the learning rate from $1\times 10^{-6}$ to the base value, followed by cosine decay to a minimum learning rate of $1\times 10^{-5}$ by the end of training. To mitigate stochastic variations, we fix the random seed to 1 throughout model training and inference.
A.6 Band Structure Results
We examine the accuracy and physical reliability of the band structures predicted by our method, by comparing the results obtained from three different Hamiltonians on representative testing samples spanning diverse elements and structures, as illustrated in Fig. 11. The red curves correspond to the ground-truth bands derived from the self-consistent Hamiltonian $\mathbf{H}^{gt}=\mathbf{H}^{(T)}$; the blue curves correspond to the bands obtained from the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$; and the orange curves represent the bands obtained from the predicted Hamiltonian of our full method, $\widehat{\mathbf{H}}=\mathbf{H}^{(0)}+\widehat{\Delta\mathbf{H}}$. The results show that the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$ provides only a rough sketch of the band structure: it approximately captures the overall positions and qualitative trends of the bands, but suffers from noticeable deviations in curvature and energy levels. In contrast, after applying neural corrections, the predicted Hamiltonian $\widehat{\mathbf{H}}$ yields band structures that align almost perfectly with the DFT ground truth, showing no significant deviations. This striking agreement demonstrates the practical value of our method for materials science and technology, where obtaining accurate band structures is a central problem.
Figure 11: Comparison of band structures obtained from Hamiltonians of representative testing samples. For each subfigure, the left and right panels show different comparisons. In both panels, the red solid curves correspond to the ground-truth bands derived from the self-consistent Hamiltonian $\mathbf{H}^{gt}=\mathbf{H}^{(T)}$. In the left panel, the black dashed curves represent the bands from the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$. In the right panel, the blue dashed curves represent the bands from the predicted Hamiltonian of our full method, $\widehat{\mathbf{H}}=\mathbf{H}^{(0)}+\widehat{\Delta\mathbf{H}}$.
A.7 Efficiency Comparison between Our Method and DFT
We evaluate the efficiency of NextHAM against the conventional DFT workflow on the same Linux server equipped with Intel(R) Xeon(R) Silver 4114 CPUs@2.20 GHz and NVIDIA A800 (80 GiB) GPUs. All DFT computations are executed on the CPU, while the neural inference of NextHAM is evaluated on both CPU and GPU. On the CPU, both DFT and NextHAM are run with four CPU cores in parallel. On the GPU, we use four A800 cards: each card executes one neural-network sub-model, and the outputs are aggregated on a single card. The testing batch size is fixed to 1 (no batching). We report the minimum, mean, and maximum wall-clock times across all testing samples.
Table 3 summarizes the runtime results:
For DFT, the entry $\mathbf{H}^{(0)}$@CPU (stage 1) includes reading the structural inputs from disk and constructing the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$ from scratch. This stage performs no diagonalization. The entry SC@CPU (stage 2) measures the self-consistent (SC) loop that starts from $\mathbf{H}^{(0)}$ and iterates to the converged $\mathbf{H}^{(T)}$ with repeated matrix diagonalizations. Writing the final results to disk is also included in this stage. The entry Total: $\mathbf{H}^{(0)}$@CPU + SC@CPU is the total runtime for the DFT workflow, i.e., the sum of the two stages.
For NextHAM, the stage 1, i.e., $\mathbf{H}^{(0)}$@CPU has the same meaning as in DFT: the cost of constructing $\mathbf{H}^{(0)}$ from the initial electron density. The stage 2, i.e., NN@CPU or NN@GPU, covers the full inference workflow after $\mathbf{H}^{(0)}$ is available: loading $\mathbf{H}^{(0)}$ into the model’s input tensors, running the neural-network forward pass to predict $\Delta\mathbf{H}$, post-processing the outputs into a DFT-compatible Hamiltonian format, and writing the results to disk. The rows Total: $\mathbf{H}^{(0)}$@CPU + NN@CPU and Total: $\mathbf{H}^{(0)}$@CPU + NN@GPU report end-to-end runtimes of NextHAM with neural inference on CPU or GPU, respectively.
From Table 3, we could observe that, NextHAM is substantially faster than the conventional DFT pipeline. Using GPU inference, the mean wall-clock time drops from 2307.11 s (DFT total) to 58.47 s (97.4% time reduction). Even with CPU inference, the mean time is 68.08 s (97.0% time reduction). In the worst case, the total runtime decreases from 28617.18 s to 744.66 s with GPU inference (97.3% time reduction), and to 755.84 s with CPU inference (97.3% speedup).
Within the DFT workflow, the self-consistent (SC) stage constitutes the overwhelming majority of the runtime, accounting for 97.6% of the mean total (2251.64 s out of 2307.11 s) and 99.2% of the observed maximum (28397.45 s out of 28617.18 s). This observation is consistent with its algorithmic structure: each SC iteration entails dense matrix diagonalizations with computational complexity $\mathcal{O}(N^{3})$, leading to an overall cost of $\mathcal{O}(TN^{3})$, where $N$ denotes the atom number in a cell and $T$ denotes the number of SC iterations. Since $T$ may be very large and is strongly problem dependent, with no reliable a priori bound on convergence, wall-clock times are both substantial and difficult to predict, and the worst-case runtime can be prohibitive. By contrast, NextHAM avoids the iterative SC loop entirely. As discussed in previous sections, constructing $\mathbf{H}^{(0)}$ scales with the number of non-zero Hamiltonian elements and is $\mathcal{O}(N^{2})$ for small systems, crossing over toward $\mathcal{O}(N)$ for sufficiently large ones; the neural inference follows the same scaling and produces a result in a single forward pass. This one-shot computation makes the runtime more predictable and markedly lower in both mean and worst-case scenarios. Moreover, neural inference benefits strongly from hardware parallelism: switching from CPU to GPU significantly reduces the mean inference time.
It is worth noting that our testing set does not contain very large systems, and the number of non-zero Hamiltonian entries typically scales as $N^{2}$ (many atoms fall within each other’s cutoff). Even in this less favorable sparsity regime, NextHAM already delivers the large speedups reported in Table 3. For substantially larger systems, the neighbor count of each atom saturates and the total number of non-zero elements grows only as $\mathcal{O}(N)$, so both $\mathbf{H}^{(0)}$ construction and neural inference become near-linear, while DFT remains $\mathcal{O}(TN^{3})$ with an a priori unknown iteration count $T$. Hence the efficiency advantage of NextHAM over DFT should increase further at scale as the system becomes larger. We point out that in our current CPU implementation the construction of $\mathbf{H}^{(0)}$ accounts for a large portion of the runtime. Fortunately, this step requires no matrix diagonalization and can be carried out in a highly parallel fashion. In future work, we plan to exploit GPU-based parallel algorithms for $\mathbf{H}^{(0)}$ preparation, which is expected to dramatically reduce this overhead and further amplify the efficiency advantage of NextHAM. We leave these works as future work plans.
Overall, the combination of favorable scaling, single-pass prediction (no SC iterations), and efficient GPU parallelization enables NextHAM to deliver large speedups across the board, opening a practical path to high-throughput materials simulations.
Table 3: Runtime on the testing set of Materials-HAM-SOC (min/max/mean seconds per sample). All stage timings include the data I/O associated with that stage. Note that the total times are computed per sample as the sum of the corresponding stages; therefore their min/max need not equal the sum of the per-stage minima/maxima.
Method Stage Min (s) Max (s) Mean (s)
DFT $\mathbf{H}^{(0)}$@CPU 3.14 742.43 55.46
SC@CPU 16.01 28397.45 2251.64
Total: $\mathbf{H}^{(0)}$@CPU + SC@CPU 21.86 28617.18 2307.11
NextHAM $\mathbf{H}^{(0)}$@CPU 3.14 742.43 55.46
NN@CPU 5.15 26.92 12.62
NN@GPU 1.16 8.95 3.01
Total: $\mathbf{H}^{(0)}$@CPU + NN@CPU 12.69 755.84 68.08
Total: $\mathbf{H}^{(0)}$@CPU + NN@GPU 4.84 744.66 58.47
A.8 Ablation Studies
We conduct fine-grained ablation studies for our framework by comparing the following settings. All ablation variants are implemented by removing a single component from the Full Method of NextHAM, while keeping all other settings identical, so as to validate the effect of each component:
Ablation@Input: In this ablation term, we replace the zeroth-step Hamiltonians in our input descriptors with conventional atom (node) and atomic-pair (edge) embeddings. Specifically, for an atom $a$ of chemical element $Z_{a}$, we maintain a learnable 32-dimensional embedding vector $\mathbf{e}_{a}=\mathbf{e}_{Z_{a}}\in\mathbb{R}^{32}$, randomly initialized and updated during network training. The embedding of an atomic pair $(a,b)$ is the concatenation of the two element embeddings, $\mathbf{e}_{ab}=[\mathbf{e}_{Z_{a}};\,\mathbf{e}_{Z_{b}}]\in\mathbb{R}^{64}$.
Ablation@Output: In this ablation term, the residual learning scheme, in which the network predicts the correction term $\Delta\mathbf{H}=\mathbf{H}^{(T)}-\mathbf{H}^{(0)}$, is removed. Instead, the neural network is trained to directly regress the full self-consistent Hamiltonian $\mathbf{H}^{(T)}$, following the setting commonly adopted in existing deep learning approaches for Hamiltonian prediction. This ablation allows us to examine the effectiveness of using $\Delta\mathbf{H}$ as the output target in reducing the complexity of the regression space and improving generalization.
Ablation@TraceGrad: In this ablation term, we remove the TraceGrad mechanism. Concretely, the supervision from the trace quantity is omitted in the loss function, and the gradient-based mechanism that delivers non-linearity from SO(3)-invariant features $z^{(\text{edge})}_{ab}$ to induce SO(3)-equivariant features via $\frac{\partial\,z^{(\text{edge})}_{ab}}{\partial\,\mathbf{f}^{\prime\text{(edge)}}_{ab}}$ is also discarded.
Ablation@Ensemble: In this ablation term, we remove the ensemble mechanism based on distance ranges. Instead of training multiple sub-models specialized for different interatomic distance intervals and aggregating their outputs, a single neural network is used to predict all Hamiltonian correction terms across all distance ranges.
Ablation@Loss-k: In this ablation term, we remove the $k$-space loss terms and train the neural network using only the real-space loss, as is commonly used in most of the existing deep learning approaches for Hamiltonian prediction. This setting allows us to assess the contribution of the $k$-space supervision in improving the physical fidelity of the predicted Hamiltonians and the resulting band structures, particularly in eliminating ghost states.
Ablation@Loss-PQ: This variant retains the $k$-space supervision on the intra-subspace blocks ($\mathcal{P}$ and $\mathcal{Q}$) but removes the cross-subspace coupling penalty, i.e., we set $\lambda_{PQ}=0$. This ablation isolates the role of the $PQ$ term.
Full Method of NextHAM.
Table 4: Comparison of Gauge MAE computed in real space (R-space) for different ablation terms and the full method on the testing set of Materials-HAM-SOC. Metrics are averaged over non-zero elements only; entries set to zero due to the truncation distance are masked out. All values are in meV.
Method Gauge MAE (meV)
Ablation@Input 1.720
Ablation@Output 2.974
Ablation@TraceGrad 1.789
Ablation@Ensemble 1.862
Ablation@Loss-k 1.615
Ablation@Loss-PQ 1.496
Full Method 1.417
We train all of the ablation terms under the same number of epochs, optimizer, and scheduler as the full method (see Appendix A.5), then evaluate them on the testing set. The $R$-space errors are summarized in Table 4. Beyond $R$-space, because $k$-space is directly tied to downstream quantities (e.g., band structures), we visualize band predictions for Ablation@Loss-k (R-space only), Ablation@Loss-PQ (setting $\lambda_{PQ}=0$) versus the Full Method in Fig. 14.
From Table 4, the Full Method achieves the lowest Gauge MAE. The Full Method reduces the error by $17.6\%$, $52.3\%$, $20.7\%$, $23.8\%$, $12.2\%$, and $5.2\%$ compared with Ablation@Input, Ablation@Output, Ablation@TraceGrad, Ablation@Ensemble, Ablation@Loss-k, and Ablation@Loss-PQ, respectively. As shown in Fig. 14, Ablation@Loss-k, which removes the $k$-space supervision and relies solely on real-space loss, produces band structures with frequent ghost states: in many cases, while most $k$-points are predicted reasonably well, some $k$-points exhibit abrupt and severe deviations from the ground truth—hallmarks of non-physical artifacts. This phenomenon mainly arises from the error amplification effect analyzed in Appendix A.2, where the large condition number of the overlap matrix can magnify small real-space errors into significant $k$-space deviations. Importantly, such sparse but catastrophic failures cannot be effectively captured by real-space loss alone. Ablation@Loss-PQ, which augments the training with $k$-space supervision on the intra-subspace blocks ($P$ and $Q$), demonstrates better performance than Ablation@Loss-k, but still fails to completely suppress ghost states. The reason is that unphysical couplings between the low-energy subspace $P$ and the high-energy subspace $Q$ remain unpenalized, and these couplings are precisely what give rise to unphysical artifacts in the band structures. In contrast, the Full Method introduces an important penalty on the $PQ$ cross block, which has clear physical significance: for the exact Hamiltonian, $P$ and $Q$ are strictly decoupled, and any spurious $PQ$ couplings in the predicted Hamiltonian are the direct source of ghost states. By explicitly enforcing this decoupling, the $PQ$ loss term addresses the root cause of the artifacts. As a result, the full method produces band structures in excellent agreement with first-principles DFT and free of ghost states. This comparison clearly demonstrates the necessity of our $k$-space loss design, in particular the $PQ$ penalty, for ensuring the physical reliability of predicted band structures.
These results collectively indicate that injecting the physically informed zeroth-step Hamiltonian as an input prior improves generalization, and predicting $\Delta\mathbf{H}=\mathbf{H}^{(T)}-\mathbf{H}^{(0)}$ reduces the effective regression space and eases optimization. They further confirm the effectiveness of the TraceGrad mechanism: supervising with the trace quantity and propagating non-linearity from invariant to equivariant features enhances representation quality. Notably, this observation aligns with the findings of Yin et al. (2025) on simpler GNN backbones, and our results demonstrate that TraceGrad remains effective within a Transformer-based framework. Moreover, the ensemble strategy, which partitions the regression space by interatomic distance and aggregates multiple specialized sub-models, yields measurable capacity gains over a single monolithic predictor, highlighting the benefit of distance-dependent specialization. In addition, $k$-space supervision provides complementary guidance that enhances physical fidelity, while explicitly penalizing the cross-subspace coupling ($PQ$) significantly suppresses band structure errors and eliminates unphysical artifacts. In summary, all validated components contribute both individually and synergistically to the overall performance of our method.
Figure 14: Comparison of band structure performance on representative testing samples. For each subfigure: in all panels, the red solid curves correspond to the ground-truth bands derived from the ground-truth self-consistent Hamiltonians. In the left panel, the black dashed curves represent the band structure results of the ablation term Ablation@Loss-k, which exhibit ghost states, i.e., abrupt and severe deviations from the ground truth at certain $k$-points. In the middle panel, the green dashed curves correspond to the results of Ablation@Loss-PQ, where such artifacts are mitigated but not fully removed. In the right panel, the blue dashed curves denote the predictions of our full method, which successfully eliminates ghost states and achieves excellent agreement with the ground truth.